# Hilbert property for low-genus families and degree-one del Pezzo surfaces

### v1

@@ Abstract.
We prove that the Hilbert property is satisfied by certain del Pezzo surfaces of degree one and Picard rank 1 over fields finitely generated over $\mathbb{Q}$. We generalize results of the first author on elliptic surfaces and employ constructions used by Desjardins and the third author to prove density of rational points. Our results are the first on the Hilbert property for minimal del Pezzo surfaces of degree one without a conic fibration.
@@ 1. Introduction
For the definitions of Hilbertian fields and the Hilbert property, see Section 2.1.
The Hilbert property takes its name from Hilbert’s Irreducibility Theorem on specializations of polynomials, which establishes the Hilbert property for $\mathbb{A}^{n}_{K}$ for $K$ a number field. The Hilbert property is closely connected to the inverse Galois problem: if all unirational varieties over a number field $K$ satisfy the Hilbert property, then for any finite group $G$, there exists a Galois extension $L/K$ with $\operatorname{Gal}(L/K)\cong G$ [24, Proof of Thm. 3.5.9].
In this paper we generalize a result of the first author in [7, Thm. 1.2] proving the Hilbert property for surfaces over a number field with multiple genus-$1$ fibrations which are sufficiently transversal (Definition 2.5). In our adaptation of this theorem, we allow one of the elliptic fibrations to be replaced by a low-genus family (Definition 2.7).
@@
Let $S$ be a smooth projective geometrically connected surface over a field $K$ which is finitely generated over $\mathbb{Q}$. Suppose given genus-$1$ fibrations $\pi_{i}:S\rightarrow\mathbb{P}^{1}$, $1\leq i\leq n$ and a low-genus family $\phi:X\rightarrow S$ with geometrically integral generic fiber. Let $F\subset S$ denote the union of divisors on $S$ which are vertical with respect to each $\pi_{i}$ and do not intersect a general elements of $\phi$ transversally. If $S\setminus F$ is algebraically simply connected and $X(K)\subset X$ is Zariski-dense, then $S$ satisfies the Hilbert property over $K$.
Recall that a smooth variety $V$ is algebraically simply connected if any finite surjective morphism $V^{\prime}\rightarrow V$ of degree at least $2$ is ramified.
@@
We apply Theorem 1.1 to certain families of del Pezzo surfaces (smooth projective varieties $X$ of dimension $2$ with ample anticanonical class $K_{X}$). It may be that all del Pezzo surfaces with a rational point are unirational and so within the remit of the above question [21, Rem. 9.4.11]. Since del Pezzo surfaces of degree one always possess a rational point (the base-point of the anticanonical linear system), we are thus motivated to prove that they satisfy the Hilbert property.
@@
The arithmetic complexity of del Pezzo surfaces is notionally governed by the degree $d=(-K_{X})^{2}\in\{1,\dots,9\}$. Let $X/k$ be a degree-$d$ del Pezzo surface over a Hilbertian field $k$ with a rational point. For $d\geq 5$, such $X$ is rational [21, Thm. 9.4.29]; as the Hilbert property is a birational invariant of smooth varieties, Hilbert’s original result then implies the Hilbert property. The Hilbert property is established for $d=4$ in work of the second author [25]. Previous joint work of the authors [9] establishes the Hilbert property when $d=2$ and $X/k$ contains a rational point outside a closed subset of $X$. This result immediately implies the Hilbert property for $d=3$ [9, Cor. 3.7].
@@
When $d=1$, the first two authors [8] proved the Hilbert property when $k$ is a number field and $X$ possesses a general conic fibration. Little is known about rational points on minimal degree-one del Pezzo surfaces with no conic fibration, which are those of Picard rank 1. Desjardins and the third author [11] proved density of rational points for an infinite family of del Pezzo surfaces of degree one. This was generalized by Nijgh [20]. Salgado and van Luijk [22] proved that the collection of $\mathbb{Q}$-rational del Pezzo surfaces with dense rational points is dense in a parameter space for real del Pezzo surfaces, following earlier work over $\mathbb{Q}$ of Várilly-Alvarado [29], Ulas [26, 27], Ulas and Togbé [28] and Jabara [15].
The primary result of this paper is the following.
@@
By [20, Thm. 3.1], the existence of a point $P$ as in Theorem 1.2 is equivalent to Zariski-density of $S(K)$ since $K$ is finitely generated over $\mathbb{Q}$.
@@
The proof of Theorem 1.2 makes use of a geometric construction, present in the aforementioned works of Desjardins and the third author [11] and Nijgh [20], which produces a low-genus family on $\mathcal{E}$ sufficiently transversal to $\pi$ to prove density of $S(K)$.
Lastly, we show that, as $a_{4}$ and $a_{6}$ and $f$ vary, the collection of $S$ as in Theorem 1.2 with non-thin rational points and Picard rank $1$ is itself non-thin.
@@ 1.1. Acknowledgements
The authors thank Daniel Loughran for useful discussions and both the University of Bath and UniDistance Suisse for their hospitality during two week-long workshops. We also thank David Loeffler for financial support during the workshop at UniDistance Suisse. This work was supported by a Focused Research Grant from the Heilbronn Institute for Mathematical Research. The second author is supported by the University of Bristol and the Heilbronn Institute for Mathematical Research. Over the course of this project, the third author was supported by UKRI Fellowship MR/T041609/2, ERC Horizon 2020 Consolidator Grant ID 101001051, and MSCA Postdoctoral Fellowship 101148712.
@@ 2.1. Hilbert property
Without loss of generality, we can (and do) assume that each $Y_{i}$ is normal and geometrically integral, and each $f_{i}$ is finite [25, Rems. 2.2, 2.3].
@@
Number fields are Hilbertian by Hilbert’s Irreducibility Theorem; more generally and crucially for us, any field finitely generated over $\mathbb{Q}$ is Hilbertian [12, §13.3]. Local fields, finite fields and algebraically closed fields are not Hilbertian.
@@ 2.2. Low-genus families
Let $\pi:Y\rightarrow B$ be a surjective flat morphism of varieties over a field $k$.
  1. (1)
    For $g\in\mathbb{Z}_{\geq 0}$, we say that $\pi$ is a genus-$g$ fibration if the generic fiber of $\pi$ is a smooth curve of genus $g$.
  2. (2)
    We say that $\pi$ is a low-genus fibration if it is a genus-$g$ fibration for $g\in\{0,1\}$.
  3. (3)
    We say that $\pi$ is an elliptic fibration if it is a genus-$1$ fibration with a section. If further $Y$ has dimension $2$, then we call $Y$ an elliptic surface over $B$.
Before defining low-genus families, we first define transversal intersection and vertical/horizontal ramification of low-genus fibrations.
@@
Let $X$ be a variety over a field $k$ and let $C$ and $D$ be two divisors on $X$. We say that $C$ and $D$ intersect transversally if the pullback of $C\cap D$ to the desingularization of $C$ contains a reduced point.
@@
Let $\pi\colon Y\rightarrow B$ be a low-genus fibration over a curve $B$ and $\psi:Z\rightarrow Y$ be a cover with branch divisor $D$. We say that $\psi$ is vertically ramified with respect to $\pi$ if $D$ is contained in a union of fibers of $\pi$ (i.e. $D$ is vertical with respect to $\pi$) and horizontally ramified otherwise.
We are now ready to define low-genus families.
Definition 2.7.
Let $X$ be a variety over a field $k$. A low-genus family on $X$ is a surjective morphism $\phi:Y\rightarrow X$ from a variety $Y$ with a low-genus fibration $\pi:Y\rightarrow B$. The elements of the family are the images $\phi(F)$ for $F$ a fiber of $\pi$.
The following result reduces the proof of Theorem 1.1 to a special case.
@@
The proof of [7, Thm. 1.1] shows that, for $K$ a number field, given covers $\psi_{j}:Z_{j}\rightarrow S$, $j=1,\dots,s$ of degrees $\geq 2$ such that $S(K)\setminus\bigcup_{j=1}^{s}\psi_{j}(Z_{j}(K))$ is not dense, there is a subset $\{j_{1},\ldots,j_{t}\}\subset\{1,\ldots,s\}$ such that $\psi_{j_{i}}:Z_{j_{i}}\rightarrow S$ is contained in $\mathcal{C}$ for all $j_{i}\in\{j_{1},\ldots,j_{t}\}$ and $S(K)\setminus\bigcup_{i=1}^{t}\psi_{j_{i}}(Z_{j_{i}}(K))$ is not dense (the simply connected hypothesis in loc. cit. ensures that $\mathcal{C}=\emptyset$ in that setting). Since this non-density would contradict our hypothesis, we deduce the result for number fields. To deduce the same for $K$ finitely generated over $\mathbb{Q}$, note that the Mordell–Weil theorem, Faltings’ theorem and Merel’s theorem, which are the key ingredients in the proof of loc. cit. in the setting of number fields, all admit generalizations to such $K$: the generalization of Mordell–Weil follows from the Lang–Néron theorem (see [5, Cor. 7.2]), the generalization of Faltings’ theorem follows from the work of McQuillan [19], and the generalization of Merel’s theorem is established by Colliot-Thélène in [11, Appendix]. ∎
@@
Let $\mathcal{C}$ be the set in Theorem 2.8, and let $\{\psi_{i}:V_{i}\rightarrow S,i=1,\dots,s\}$ be a finite subset of $\mathcal{C}$. We will show that $S(K)\setminus\bigcup_{i=1}^{r}\psi_{i}(V_{i}(K))$ is Zariski-dense in $S$, so that we can deduce the result from Theorem 2.8. For each $i\in\{1,\dots,s\}$, denote by $F_{i}$ the fiber product $X\times_{\phi,\psi_{i}}V_{i}$. Let $X_{\eta}$ (resp. $F_{i,\eta}$) be the generic fiber of the low-genus fibration $X\rightarrow B$ (resp. $F_{i}\to X\rightarrow B$). We first show that for all $i\in\{1,\dots,s\}$, the curve $F_{i,\eta}$ is integral over $K(B)$. We have the following commutative diagram.
@@
Since $S$ is a smooth surface and $V_{i}$ is normal, the morphism $\psi_{i}$ is flat by miracle flatness [1, Tag 00R4]. It follows that $F_{i,\eta}\to X_{\eta}$ is finite flat as well (both finiteness and flatness being preserved under base change), with generic fiber $\operatorname{Spec}K(X)\times_{\operatorname{Spec}K(S)}\operatorname{Spec}K(V_{i})\cong\operatorname{Spec}(K(X)\otimes_{K(S)}K(V_{i}))$. Since $\phi$ has geometrically integral generic fiber, $K(X)\otimes_{K(S)}K(V_{i})$ is a field: the extensions $K(X)/K(S)$ and $K(V_{i})/K(S)$ are regular and algebraic respectively, hence linearly disjoint. Thus the flat morphism $F_{i,\eta}\to X_{\eta}$ has integral generic fiber, and thus $F_{i,\eta}$ is also integral [17, Prop. 4.3.8].
For each $i\in\{1,\ldots,s\}$, denote by $D_{i}\subset S$ the branch locus of $\psi_{i}$, which is a divisor by Nagata–Zariski purity [1, Tag 0BMB]. We claim that a general element of the low-genus family intersects each $D_{i}$ transversally: otherwise, there is at least one $i_{0}$ such that $D_{i_{0}}$ belongs to $F$, but then the restriction of $\psi_{i_{0}}$ over $S\setminus F$ is an unramified cover of degree $\geq 2$, contradicting that $S\setminus F$ is algebraically simply connected. Then there is at least one smooth point on $F_{i,\eta}$ which is ramified over $X_{\eta}$ (cf. [6, p. 595]), hence $F_{i,\eta}$ is an integral $K(B)$-curve ramified over $X_{\eta}$. We may spread out over some open $U\subset B$ to a morphism $\xi_{i}\colon F_{i,U}\to X_{U}$ such that, for all $P\in U$, the curve $F_{i,P}$ is a ramified cover of $X_{P}$.
Since $K$ is finitely generated over $\mathbb{Q}$ and $X(K)$ is Zariski-dense, there is a non-empty Zariski open $V\subset X_{U}$ such that, for each $v\in V(K)$, the fiber $\pi^{-1}(\pi(v))$ has infinitely many $K$-rational points; see [11, Appendix] for the genus-$1$ case (the genus-0 case is elementary, as any smooth curve of genus 0 with one rational point has infinitely many). By assumption, $V(K)$ is Zariski-dense in $V$, and we deduce Zariski-density in $U$ of

### v2

@@ Abstract.
We prove that the Hilbert property is satisfied by certain del Pezzo surfaces of degree one and Picard rank one over fields finitely generated over $\mathbb{Q}$. We generalize results of the first author on elliptic surfaces and employ constructions used by Desjardins and the third author to prove density of rational points. Our results are the first on the Hilbert property for minimal del Pezzo surfaces of degree one without a conic fibration.
@@ 1. Introduction
Informally, the Hilbert property holds for a variety $X$ defined over a field $K$ if Hilbert’s Irreducibility Theorem (from which it takes its name) holds for polynomials with coefficients in the function field $K(X)$, and it is a measure of the abundance of rational points of $X$. Hilbert’s original theorem essentially establishes the Hilbert property for $\mathbb{A}^{n}_{K}$ for $K$ a number field. See Section 2.1 for the precise definition. The Hilbert property is closely connected to the inverse Galois problem: if all unirational varieties over a number field $K$ satisfy the Hilbert property, then for any finite group $G$, there exists a Galois extension $L/K$ with $\operatorname{Gal}(L/K)\cong G$ [28, Proof of Thm. 3.5.9].
A now well-developed technique to prove the Hilbert property for a surface $X$ is to produce an “abundance” of curves of geometric genus 0 or $1$ on $X$ and use these to propagate rational points. When one is able to produce an abundance of curves of genus 0 (e.g. at least enough to infer the unirationality of $X$), one usually hopes to prove more than the Hilbert property: see for instance the paper [30] by Swinnerton–Dyer and the papers [10, 11] by the authors of this paper, where weak weak approximation of the surfaces under consideration is proven. When one is able to produce many curves of genus $1$, but not (necessarily) curves of genus 0, the Hilbert property is usually considered the best hope with the current technology.
Historically, the first use of curves of genus $1$ in this context seems to be Manin’s proof of the Hilbert property of cubic surfaces with a rational point [22, §46], by employing the $2$-dimensional linear system of cubic curves on the surface. (Consistently with the philosophy laid out above, the aforementioned result of Swinnerton–Dyer strenghtened the result of Manin by ingeniously using the curves of genus 0 as well.) After Manin, not much work was done in this direction until the paper of Corvaja and Zannier [8] proving the Hilbert property for the K3 surface $x^{4}+y^{4}=z^{4}+w^{4}$ using two elliptic fibrations. Their result was generalized by the first-named author of this paper to any surface with multiple genus-$1$ fibrations and Zariski-dense rational points, assuming a simple connectedness hypothesis for a certain Zariski-open subset of the surface [9, Thm. 1.2]. Since then, the multiple fibrations method has appeared in multiple works in analogous contexts: see for instance [29] (for conic bundles), [16, 17] (for the potential analogue), and [4, 5, 1] (for the integral analogue).
In this paper, we generalize [9, Thm. 1.2] by replacing one of the fibrations with a general family of curves of low genus (0 or 1). Unlike in Manin’s work and in the multiple fibrations method, our family may be non-linear (as it is in our main application, Theorem 1.2). This introduces an additional difficulty: in the linear case, every rational point lies on at least one member of the family defined over $K$, and this is no longer true in the non-linear setting. A similar use of non-linear families to study the Hilbert property first appeared in the work [5] of Coccia in the analogous context of integral points, but to our knowledge this paper is the first to apply this idea to rational points and to establish a general result of the following kind.
@@
Let $S$ be a smooth projective geometrically connected surface over a field $K$ which is finitely generated over $\mathbb{Q}$. Suppose we are given $n\geq 0$ genus-11 fibrations $\pi_{i}:S\rightarrow\mathbb{P}^{1}$, $1\leq i\leq n$ and a low-genus family of curves $\{X_{b}\}_{b\in B}$ on $S$ with some base $B/K$ such that the induced map $\phi\colon X\to S$ from their total space $X$ is dominant with geometrically integral generic fiber. Let $F\subset S$ denote the union of divisors on $S$ which are vertical with respect to each $\pi_{i}$ and do not intersect a general member of the family $\{X_{b}\}_{b\in B}$ transversally. If $S\setminus F$ is algebraically simply connected and $X(K)\subset X$ is Zariski-dense, then $S$ satisfies the Hilbert property over $K$.
Recall that a smooth variety $V$ is algebraically simply connected if any finite surjective morphism $V^{\prime}\rightarrow V$ of degree at least $2$ is ramified. The assumption that $\phi$ has geometrically integral generic fiber is necessary for the method we employ to work (see Remark 2.8). The other hypothesis of Zariski-density on $X$, standing in place of the more common and natural hypothesis of Zariski-density of rational points on $S$, arises from the “non-linear difficulty” mentioned above.
@@
We apply Theorem 1.1 to certain families of del Pezzo surfaces (smooth projective varieties $X$ of dimension $2$ with ample anticanonical class $K_{X}$). It may be that all del Pezzo surfaces with a rational point are unirational and so within the remit of the above question [25, Rem. 9.4.11]. Since del Pezzo surfaces of degree one always possess a rational point (the base-point of the anticanonical linear system), we are thus motivated to prove that they satisfy the Hilbert property.
@@
The arithmetic complexity of del Pezzo surfaces is notionally governed by the degree $d=(-K_{X})^{2}\in\{1,\dots,9\}$. Let $X/k$ be a degree-$d$ del Pezzo surface over a Hilbertian field $k$ with a rational point. For $d\geq 5$, such $X$ is rational [25, Thm. 9.4.29]; as the Hilbert property is a birational invariant of smooth varieties, Hilbert’s original result then implies the Hilbert property. The Hilbert property is established for $d=4$ in work of the second author [29]. Previous joint work of the authors [11] establishes the Hilbert property when $d=2$ and $X/k$ contains a rational point outside a closed subset of $X$. This result immediately implies the Hilbert property for $d=3$ [11, Cor. 3.7].
@@
When $d=1$, the first two authors [10] proved the Hilbert property when $k$ is a number field and $X$ possesses a general conic fibration. Little is known about rational points on minimal degree-one del Pezzo surfaces with no conic fibration, which are those of Picard rank 1. Desjardins and the third author [13] proved density of rational points for an infinite family of del Pezzo surfaces of degree one. The surfaces in this family are endowed with a family of low-genus curves contained (non-linearly) in $|{-3K_{S}}|$. It is this family of curves that Desjardins and the third author use to prove density. Nijgh [24] later generalized the work [13] by enlarging the family of surfaces to which the method can be applied. Salgado and van Luijk [26] proved that $\mathbb{Q}$-rational del Pezzo surfaces with dense rational points are dense in a parameter space for real del Pezzo surfaces, following work of Várilly-Alvarado [34], Ulas [31, 32], Ulas and Togbé [33] and Jabara [19].
The primary result of this paper is the proof of the Hilbert property for the family of surfaces for which Nijgh proves Zariski-density of rational points.
@@
By [24, Thm. 3.1], the existence of a point $P$ as in Theorem 1.2 is equivalent to Zariski-density of $S(K)$ since $K$ is finitely generated over $\mathbb{Q}$.
@@
We prove Theorem 1.2 by applying Theorem 1.1 to the surface $\mathcal{E}$ endowed with one elliptic fibration $\pi$ and with a certain low-genus family of curves introduced in the aforementioned works of Desjardins and the third author [13] and generalized by Nijgh [24], whose construction we sketch in the next paragraph.
The del Pezzo surfaces appearing in the family of Theorem 1.2 are endowed with a 33-to-11 morphism $\theta:S\to W$ onto a (singular) cubic surface $W$ in $\mathbb{P}^{3}_{K}$. Given a rational point $P\in S(K)$ lying in the inverse image $S^{o}:=\theta^{-1}(W^{o})$ of the smooth locus $W^{o}\subset W$, one may consider the intersection curve $D_{P}:=T_{\theta(P)}W\cap W$. For a general $P$, this is a nodal curve in $P$ with geometric genus zero. The arithmetic genus of its inverse image $\theta^{-1}(D_{P})\in|{-3K_{S}}|$ is easily computed to be $4$ via the adjunction formula. Moreover, the curve $C_{P}:=\theta^{-1}(D_{P})$ has three nodal singularities at the (geometric) points $\{P,Q_{1},Q_{2}\}=\theta^{-1}(\theta(P))$, making its geometric genus at most 43=14-3=1. The low-genus family is then the associated family $\{\rho^{-1}(C_{P})\}_{P\in S^{o}}$ in $\mathcal{E}$.
As presented in [24] (following the idea of [13]), this family may already be directly used to obtain Zariski-density of rational points on $S$. For each $p\in\{P,Q_{1},Q_{2}\}$, the point [2]p[-2]p also lies on $C_{P}$, where $2]:S\dashrightarrow S$ denotes the rational map induced by the self-isogeny [2][-2] of $\mathcal{E}$. The zero-cycle $2](Q_{1})+[-2](Q_{2})-2([-2]P)\in Z_{0}(S\otimes_{K}\bar{K})$ then lies on $C_{P}$ and is $\operatorname{Gal}(\bar{K}/K)$-invariant, and thus defines a rational point in the Jacobian of (the normalization of) $C_{P}$. Following [13] and [24], this element is non-torsion for $P$ lying outside of a thin set, and thus produces infinitely many rational points on $C_{P}$.
The just-sketched argument for Zariski-density proves that the total space $\mathcal{C}$ of the family $\{C_{P}\}_{P\in S^{o}}$ has Zariski-dense rational points, verifying one of the hypotheses of Theorem 1.1. The other two hypotheses are verfied in Propositions 3.7 and 3.12: the first proves that the union $F$ is empty in our case, and the second proves that $\phi:C\to S$ has geometrically integral generic fiber.
Lastly, we show that, as $a_{4}$, $a_{6}$ and $f$ vary, the collection of $S$ as in Theorem 1.2 with non-thin rational points and Picard rank $1$ is itself non-thin.
@@ 1.1. Acknowledgements
The authors thank Daniel Loughran for useful discussions and both the University of Bath and UniDistance Suisse for their hospitality during two week-long workshops. We also thank David Loeffler for financial support during the workshop at UniDistance Suisse. Lastly we thank Jean-Louis Colliot-Thélène for drawing our attention to the work of Manin [22]. This work was supported by a Focused Research Grant from the Heilbronn Institute for Mathematical Research. The second author is supported by the University of Bristol and the Heilbronn Institute for Mathematical Research. Over the course of this project, the third author was supported by UKRI Fellowship MR/T041609/2, ERC Horizon 2020 Consolidator Grant ID 101001051, and MSCA Postdoctoral Fellowship 101148712.
@@ 2.1. Hilbert property
Without loss of generality, we can (and do) assume that each $Y_{i}$ is normal and geometrically integral, and each $f_{i}$ is finite [29, Rems. 2.2, 2.3].
@@
Number fields are Hilbertian by Hilbert’s Irreducibility Theorem; more generally and crucially for us, any field finitely generated over $\mathbb{Q}$ is Hilbertian [15, §13.3]. Local fields, finite fields and algebraically closed fields are not Hilbertian.
@@ 2.2. Low-genus families
Let $\pi:Y\rightarrow B$ be a surjective flat morphism of varieties over a field $k$. For $g\in\mathbb{Z}_{\geq 0}$, we say that $\pi$ is a genus-$g$ fibration if the generic fiber of $\pi$ is a geometrically integral curve of geometric genus $g$. We say that $\pi$ is an elliptic fibration if it is a genus-$1$ fibration with smooth generic fiber and a section. If further $Y$ has dimension $2$, then we call $Y$ an elliptic surface over $B$.
We recall that two curves $C$ and $D$ on a smooth surface $S$ are said to intersect transversally if they intersect transversally at all points, or equivalently, if the intersection scheme $C\cap D$ is reduced. In fact, for Theorem 1.1 we only need the weaker requirement of one reduced point in the intersection.
@@
Let $\pi\colon Y\rightarrow B$ be a genus-$g$ fibration over a base $B$ with $Y$ regular, and $\psi:Z\rightarrow Y$ be a cover with branch divisor $D$. We say that $\psi$ is vertically ramified with respect to $\pi$ if $D$ is contained in a union of fibers of $\pi$ (i.e. $D$ is vertical with respect to $\pi$), and horizontally ramified otherwise.
A flat family of proper subschemes of a variety $Y/k$ with base $B/k$ is a closed subscheme $X\subset Y\times_{k}B$, flat over $B$. For such a family, we call $B$ the base of the family, $Y$ the total space of the family. Write $p$ for the projection of $Y\times_{k}B$ to $B$, and $\phi$ for the projection to $Y$. Then the subschemes $X_{b}=\phi(p^{-1}(b)),b\in B$ the members of the family. We denote such a family by $\{X_{b}\}_{b\in B}$.
@@
Let $Y$ be a proper variety over a field $k$. A flat family of proper subschemes $\{X_{b}\}_{b\in B}$ of $Y$ is a low-genus family if its generic element is a geometrically integral curve of geometric genus 0 or $1$.
With the following result, essentially contained in [9], we reduce the proof of Theorem 1.1 to a special case.
@@
The proof of [9, Thm. 1.1] shows that, for $K$ a number field, given covers $\psi_{j}:Z_{j}\rightarrow S$, $j=1,\dots,s$ of degrees $\geq 2$ such that $S(K)\setminus\bigcup_{j=1}^{s}\psi_{j}(Z_{j}(K))$ is not dense, there is a subset $\{j_{1},\ldots,j_{t}\}\subset\{1,\ldots,s\}$ such that $\psi_{j_{i}}:Z_{j_{i}}\rightarrow S$ is contained in $\mathcal{C}$ for all $j_{i}\in\{j_{1},\ldots,j_{t}\}$ and $S(K)\setminus\bigcup_{i=1}^{t}\psi_{j_{i}}(Z_{j_{i}}(K))$ is not dense (the simply connected hypothesis in loc. cit. ensures that $\mathcal{C}=\emptyset$ in that setting). Since this non-density would contradict our hypothesis, we deduce the result for number fields. To deduce the same for $K$ finitely generated over $\mathbb{Q}$, note that the Mordell–Weil theorem, Faltings’ theorem and Merel’s theorem, which are the key ingredients in the proof of loc. cit. in the setting of number fields, all admit generalizations to such $K$: the generalization of Mordell–Weil follows from the Lang–Néron theorem (see [7, Cor. 7.2]), the generalization of Faltings’ theorem follows from the work of Faltings–Wüstholz [14, Thm. 3, p. 205], and the generalization of Merel’s theorem is established by Colliot-Thélène in [13, Appendix]. ∎
@@
Define $\mathcal{C}$ to be the set in Theorem 2.7 if $n\geq 1$ or the set of all finite covers of $S$ of degree $\geq 2$ if $n=0$, and let $\{\psi_{i}:V_{i}\rightarrow S,i=1,\dots,s\}$ be a finite subset of $\mathcal{C}$. We will show that $S(K)\setminus\bigcup_{i=1}^{r}\psi_{i}(V_{i}(K))$ is Zariski-dense in $S$, so that we can deduce the result from Theorem 2.7. For each $i\in\{1,\dots,s\}$, denote by $F_{i}$ the fiber product $X\times_{\phi,\psi_{i}}V_{i}$. Let $X_{\eta}$ (resp. $F_{i,\eta}$) be the generic fiber of the low-genus fibration $\pi:X\rightarrow B$ (resp. $F_{i}\to X\rightarrow B$).
Our assumption that the generic fiber of $\phi$ is geometrically integral implies that for all $i\in\{1,\dots,s\}$, the curve $F_{i,\eta}$ is integral over $K(B)$. In fact, consider the following commutative diagram.
@@
Since $S$ is a smooth surface and $V_{i}$ is normal, the morphism $\psi_{i}$ is flat by miracle flatness [2, Tag 00R4]. It follows that $F_{i,\eta}\to X_{\eta}$ is finite flat as well (both finiteness and flatness being preserved under base change), with generic fiber $\operatorname{Spec}K(X)\times_{\operatorname{Spec}K(S)}\operatorname{Spec}K(V_{i})\cong\operatorname{Spec}(K(X)\otimes_{K(S)}K(V_{i}))$. Since $\phi$ has geometrically integral generic fiber, the extension $K(X)/K(S)$ is regular and thus linearly disjoint from the algebraic extension $K(V_{i})/K(S)$. Hence $K(X)\otimes_{K(S)}K(V_{i})$ is a field. In particular, the flat morphism $F_{i,\eta}\to X_{\eta}$ has integral generic fiber, and thus $F_{i,\eta}$ is also integral [21, Prop. 4.3.8].
For each $i\in\{1,\ldots,s\}$, denote by $D_{i}\subset S$ the branch locus of $\psi_{i}$, which is a divisor by Nagata–Zariski purity [2, Tag 0BMB]. We claim that a general member of the low-genus family intersects each $D_{i}$ transversally: otherwise, there is at least one $i_{0}$ such that $D_{i_{0}}$ belongs to $F$, but then the restriction of $\psi_{i_{0}}$ over $S\setminus F$ is an unramified cover of degree $\geq 2$, contradicting that $S\setminus F$ is algebraically simply connected. Then there is at least one smooth point on $F_{i,\eta}$ which is ramified over $X_{\eta}$ (cf. [8, p. 595]), hence $F_{i,\eta}$ is an integral $K(B)$-curve ramified over $X_{\eta}$. We may spread out over some open $U\subset B$ to a morphism $\xi_{i}\colon F_{i,U}\to X_{U}$ such that, for all $P\in U$, the curve $F_{i,P}$ is a ramified cover of $X_{P}$.
Since $K$ is finitely generated over $\mathbb{Q}$ and $X(K)$ is Zariski-dense, there is a non-empty Zariski open $V\subset X_{U}$ such that, for each $v\in V(K)$, the fiber $\pi^{-1}(\pi(v))$ has infinitely many $K$-rational points; see [13, Appendix] for the genus-$1$ case (the genus-0 case is elementary, as any smooth curve of genus 0 with one rational point has infinitely many). By assumption, $V(K)$ is Zariski-dense in $V$, and we deduce Zariski-density in $U$ of
@@
In the following remark we argue on the necessity of the hypothesis of geometric integrality of the generic fiber of $\phi$.
Remark 2.8.
We make some observations on the applicability of Theorem 1.1.
  1. (1)
    A special case of Theorem 1.1 is that with no genus-$1$ fibrations ($n=0$). In this case, our proof shows that in fact $\phi(X(K))$ is non-thin in $S(K)$.
  2. (2)
    We give an example of a surface $S$ and a family of genus-0 curves $\phi:X\to S$ with fibration $X\to B$ whose general member intersects any given divisor $D$ on $S$ transversally, and yet $\phi(X(K))$ is thin in $S(K)$. In this example, the generic fiber of $\phi$ is not geometrically integral, and so Theorem 1.1 (and point (1) of this remark) is not applicable.
    Take a general cubic pencil on $\mathbb{P}^{2}_{K}$, let $T$ be its base locus ($T$ consists of $9$ geometric points), and enlarge $K$ so that at least one of the points of $T$ is rational, call it $t$. Let $S$ be the blow-up of $\mathbb{P}^{2}_{K}$ in $T$, and let $E$ be the exceptional divisor above $t$. The proper transform of the cubic pencil makes $S$ into an elliptic surface with zero-section $E$. Let now $B$ be a non-empty Zariski-open of the dual projective space $(\mathbb{P}^{2}_{K})^{\vee}$ containing none of the lines passing through the (geometric) points of $T$. Let $\{l_{b}\}_{b\in B}$ be the family parametrizing the proper transforms in $S$ of the corresponding lines, and let $\{X_{b}:=[2]l_{b}\}_{b\in B}$ be the image of this family under the rational map $2]:S\dashrightarrow S$ (we restrict $B$ so that $l_{b}$ does not pass through any of the finitely many points not contained in the domain of $[2]$). Denote by $X$ the total space of the family $\{X_{b}\}_{b\in B}$. It is clear that $\phi(X(K))\subset[2](S(K))$ is thin. As we verify in the next paragraph, the general member $X_{b}$ intersects any given divisor $D$ on $S$ transversally, and thus this family provides an example as desired.
    Assume without loss of generality that $D$ is integral. The curve $X_{b}=[2]l_{b}$ intersects $D$ transversally if and only if $l_{b}$ intersects $2]^{*}D$ transversally and $l_{b}$ contains no pair of distinct points $q_{1},q_{2}\in S(\overline{K})$ with $2](q_{1})=[2](q_{2})\in D$. The general line $l_{b}\in(\mathbb{P}^{2}_{K})^{\vee}$ satisfies both these conditions: indeed the map $2]:S\dashrightarrow S$ is unramified on its domain of definition, so the divisor $2]^{*}D$ is reduced, and thus the general line intersects it transversally, while the second condition describes a one-dimensional set of lines.
Remark 2.9.
Theorem 1.1 still holds if one replaces $F$ with the set of divisors that are vertical with respect to each $\pi_{i}$ and do not intersect a general member of the family $\{X_{b}\}_{b\in B}$ transversally in at least one point. Indeed, observe that in the proof of Theorem 1.1 we did not need the general member of the family $\{X_{b}\}_{b\in B}$ to intersect the divisors in $F$ transversally at all intersections points, but it was rather enough that it intersects these divisors transversally in at least one point.